OpenStudy (anonymous):

Elosia and three friends went to the fair. Two of the friends played 4 more games than elosia. the third friend played three times as many games as elosia. Altogether they played a total of 56 games. What is the equation and the total number each person played?

OpenStudy (bahrom7893):

Let the number of games played by elosa be equal to x; then two friends played 4 more games than elosa, this means: 2*(x+4) Third played three times as many: 3*x; all together they palyed 56 games.

OpenStudy (bahrom7893):

x + 2*(x+4) + 3*x = 56 The number of games of elos + 2 ofher friends + 3rd friend

OpenStudy (bahrom7893):

x + 2x + 8 + 3x = 56 3x + 3x = 56 - 8 6x = 48 x = 8

OpenStudy (bahrom7893):

Elosa played 8 games; Her other two friends played 2*(x+4) = 2*(8+4) = 2*12 = 24 total or 12 games each; Third friend played 3*x = 3*8 = 24 games